# MS-Mix: Unveiling the Power of Mixup for Multimodal Sentiment Analysis

### v1

@@ 4.1. Datasets
The MOSI (Zadeh et al., 2016) dataset is an English benchmark for emotion recognition, consisting of 2,199 opinion segments from 93 YouTube movie reviews. It offers aligned text (transcripts), audio, and visual (facial) data, annotated for sentiment intensity on a continuous scale from -3 (strongly negative) to +3 (strongly positive) by five independent annotators, with final labels calculated from the average scores. The dataset reflects real-world complexity, including modality incongruity (e.g., sarcasm), and contains about 2.2 hours of video from 89 speakers. Widely utilized in multimodal fusion research, this dataset encompasses pre-extracted features (e.g., acoustic data, facial action units) alongside speaker metadata.
@@
The MOSEI (Zadeh et al., 2018b) dataset includes 23,453 video segments from more than 1,000 different speakers covering over 250 topics. It is the largest multimodal benchmark for sentiment and emotion analysis. The dataset provides aligned text transcripts, audio signals, and visual frames, all annotated by trained annotators with continuous sentiment intensity scores from -3 to 3 like MOSI, and six emotion intensities (happiness, sadness, anger, fear, disgust, surprise). It captures real-world communication nuances, including cross-modal conflicts, and supports research in multimodal fusion, emotion recognition, and sarcasm detection. Preprocessed features of MOSI and MOSEI are available through the CMU-Multimodal SDK (Zadeh et al., 2018c).
@@
The SIMS (Yu et al., 2020) dataset is a Chinese multimodal benchmark featuring 2,281 video segments from 200 diverse online videos (vlogs, reviews, dialogues), designed to address real-world modality asynchrony. It provides aligned text transcripts, audio, and visual data with annotations of unified sentiment scores on a continuous scale from -1 (strongly negative) to +1 (strongly positive). As the first Chinese dataset emphasizing cross-modal dynamics, SIMS supports research in asynchronous multimodal fusion, cross-modal interaction modeling, and Mandarin sentiment analysis, with raw videos and preprocessed features publicly available.
4.2. Backbones and Baseline Methods
To ensure a comprehensive comparison, we evaluated the proposed method against three mixup-based methods: Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), and $\mathcal{P}ow$Mix (Georgiou et al., 2024) on six MSA architectures: TFN (Zadeh et al., 2017), LMF (Liu and Shen, 2018), MuIT (Tsai et al., 2019), MISA (Hazarika et al., 2020), ALMT (Zhang et al., 2023), and GLoMo (Zhuang et al., 2024). These architectures represent a diverse spectrum of feature extraction architectures and fusion strategies, covering both early and recent advances in the field, as summarized in Table 1.
Table 1. Description of the six MSA backbones used in the paper
Model TFN (Zadeh et al., 2017) LMF (Liu and Shen, 2018) MuIT (Tsai et al., 2019) MISA (Hazarika et al., 2020) ALMT (Zhang et al., 2023) GLoMo (Zhuang et al., 2024)
Main encoder LSTM LSTM CNN LSTM Transformer Transformer
Fusion method Tensor fusion Low-rank Fusion Cross-attention Self-attention Cross-attention Self-attention
Published In, Year EMNLP, 2017 ACL, 2018 ACL, 2019 ACM MM, 2020 EMNLP, 2023 ACM MM, 2024
4.3. Experimental Settings
@@
We adopt the evaluation metrics established in the M-SENA framework (Mao et al., 2022) unified framework to ensure consistent and comparable results. These include binary accuracy ($ACC_{2}$) and F1-score, multiclass accuracy ($ACC_{5}$, $ACC_{7}$) and Mean Absolute Error (MAE). Among these, MAE is a regression metric. Classification accuracy values are derived by mapping the regression scores to discrete sentiment categories. The prefix "w-" denotes without neutral-labeled samples. Furthermore, to facilitate intuitive and comprehensive comparison across metrics, we have separately presented the average values (Avg.) of all the classification metrics (ACC and F1-score).
@@
All experiments were conducted on a high-performance server equipped with an NVIDIA GeForce RTX 3090 Ti GPU. The code was implemented in Python 3.10.18 using the PyTorch 2.7.1 framework with CUDA 12.6 support. For a fair comparison, all models were implemented using publicly available code, and their original experimental settings were retained. Most of the models were evaluated within the M-SENA framework (Mao et al., 2022), ensuring that all the comparison methods are conducted under consistent experimental conditions. In MS-Mix, the base mixing ratio $\lambda_{base}$ is sampled from a symmetric Beta distribution with $\alpha=2.0$ and the similarity threshold $\delta$ is fixed at 0.2. The loss weights $\xi_{1}$ and $\xi_{2}$ are set to 0.7 and 0.5, respectively. Additionally, the number of attention heads $h$ = 4 for the MOSI and MOSEI datasets, and $h$ = 6 for the SIMS dataset.
Table 2. Results of Various Approaches on The MOSI Dataset. Bold: Best performance. Underline: Second-best performance. $\dagger$}: result reported in (Mao et al., 2022). *: result reported in (Georgiou et al., 2024). "-": result was not reported in the original paper.
MODEL w-ACC2 (%) ACC2(%) w-F1-score(%) F1-score(%) ACC5(%) ACC7(%) MAE $\downarrow$ Avg.
@@
+ Manifold Mix 79.27 78.13 79.25 78.06 40.96 35.71 0.927 65.23
+ MultiMix 79.57 78.43 79.60 78.40 39.07 34.69 0.922 64.96
+ $\mathcal{P}ow$Mix 79.27 78.13 77.96 79.15 41.11 36.73 0.919 65.39
+ MS-Mix (ours) 80.18 78.72 80.05 78.52 41.25 34.84 0.904 65.59
@@
+ Manifold Mix 79.42 77.84 79.50 77.86 38.92 35.13 0.948 64.78
+ MultiMix 79.88 78.72 79.96 78.74 38.80 36.32 0.907 65.40
+ $\mathcal{P}ow$Mix 80.49 79.01 80.54 79.00 43.73 38.48 0.895 66.88
+ MS-Mix (ours) 82.16 80.90 82.12 80.81 42.42 36.88 0.893 67.55
@@
+ Manifold Mix 81.55 79.88 81.48 79.74 43.11 36.55 0.838 67.05
+ MultiMix 81.64 79.01 81.66 79.96 43.38 36.57 0.859 67.04
+ $\mathcal{P}ow$Mix* 81.01 - 80.99 - 40.65 35.00 0.904 -
+ $\mathcal{P}ow$Mix 81.11 79.65 80.64 79.41 41.79 35.25 0.880 66.31
+ MS-Mix (ours) 81.84 79.86 81.80 79.87 43.48 37.51 0.827 67.39
@@
+ Manifold Mix 83.69 82.22 83.69 82.17 46.50 41.11 0.765 69.90
+ MultiMix 83.38 81.34 83.28 81.16 47.38 41.69 0.749 69.71
+ $\mathcal{P}ow$Mix* 83.49 - 83.50 - 48.02 42.65 0.761 -
+ $\mathcal{P}ow$Mix 83.27 81.33 83.41 81.42 47.91 41.87 0.763 69.87
+ MS-Mix (ours) 83.99 82.36 83.96 82.28 48.25 41.98 0.746 70.47
ALMT 84.60 82.65 84.58 82.57 51.75 46.21 0.723 72.06
+ Manifold Mix 83.99 82.22 83.98 82.14 52.62 46.65 0.742 71.93
+ MultiMix 83.69 82.36 83.68 82.30 49.27 44.02 0.751 70.08
+ $\mathcal{P}ow$Mix* 84.30 - 84.30 - 49.24 44.25 0.741 -
+ $\mathcal{P}ow$Mix 84.62 81.63 84.68 81.65 51.96 46.27 0.726 71.80
+ MS-Mix (ours) 84.99 83.22 84.96 82.73 52.19 47.06 0.720 72.36
GLoMo 84.89 83.50 84.92 83.94 53.87 46.13 0.749 72.88
+ Manifold Mix 85.14 83.18 85.07 83.48 53.93 52.27 0.729 73.85
+ MultiMix 84.70 83.42 84.60 83.64 53.82 52.19 0.743 73.73
+ $\mathcal{P}ow$Mix 84.77 84.04 84.71 83.90 53.90 47.53 0.733 73.14
+ MS-Mix (ours) 85.66 83.55 85.58 83.82 54.58 52.90 0.724 74.35
Fig. 3. The t-SNE visualization of the original features and mixed features generated by MS-Mix (a) and $\mathcal{P}$owMix (b) on the MOSEI dataset using the MISA model. We employ a color scheme (blue/red) to differentiate the positive and negative categories, and use transparency to distinguish the original features from the mixed ones. :
4.4. Verification Performance of MS-Mix
This chapter evaluates our proposed method, MS-Mix, on three publicly available MSA datasets: MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018b), and SIMS (Yu et al., 2020). Experiments were conducted using six backbone models, representing both classical (Zadeh et al., 2017; Liu and Shen, 2018; Tsai et al., 2019; Hazarika et al., 2020) and state-of-the-art approaches (Zhang et al., 2023; Zhuang et al., 2024) in terms of model architectures and fusion strategies, to demonstrate the general applicability of MS-Mix across different frameworks. For each model, we compared against classical Mixup variants: Manifold Mixup (Manifold Mix) (Verma et al., 2019) and MultiMix (Venkataramanan et al., 2023), as well as a recently proposed mixup-based method in MSA, $\mathcal{P}ow$Mix (Georgiou et al., 2024). The comparative results are summarized in Tables 2,3,4.
Table 3. Results of Various Approaches on the MOSEI Dataset. Bold: Best performance. Underline: Second-best performance. $\dagger$}: result reported in (Mao et al., 2022). *: result reported in (Georgiou et al., 2024). "-": result was not reported in the original paper.
@@ 4.4. Verification Performance of MS-Mix
+ Manifold Mix 84.31 82.91 84.08 83.03 52.26 51.02 0.572 72.94
+ MultiMix 82.94 80.47 82.80 80.82 53.51 52.39 0.568 72.16
+ $\mathcal{P}ow$Mix 83.46 81.61 83.31 81.87 53.36 51.79 0.562 72.57
+ MS-Mix (ours) 83.63 82.40 83.23 82.36 54.00 52.48 0.557 73.02
@@
+ Manifold Mix 83.74 78.71 83.85 78.98 53.53 52.33 0.563 71.86
+ MultiMix 84.51 81.07 84.44 81.52 54.15 52.65 0.553 73.06
+ $\mathcal{P}ow$Mix 83.38 80.30 83.33 80.79 54.30 53.10 0.559 72.53
+ MS-Mix (ours) 84.65 83.30 84.28 83.26 54.13 52.74 0.554 73.73
@@
+ Manifold Mix 84.46 81.64 84.54 81.40 54.02 52.51 0.561 73.10
+ MultiMix 84.82 81.31 84.61 81.03 54.79 52.07 0.562 73.11
+ $\mathcal{P}ow$Mix* 84.44 - 84.38 - 54.26 52.75 0.559 -
+ $\mathcal{P}ow$Mix 84.37 81.45 84.56 81.34 54.57 52.45 0.560 73.12
+ MS-Mix (ours) 84.68 81.69 84.64 82.07 54.98 53.09 0.556 73.53
@@
+ Manifold Mix 84.67 80.94 84.67 81.48 53.85 52.05 0.547 72.94
+ MultiMix 84.78 80.49 84.79 81.08 53.68 52.29 0.544 72.85
+ $\mathcal{P}ow$Mix* 84.97 - 84.86 - 54.52 53.00 0.543 -
+ $\mathcal{P}ow$Mix 84.50 81.34 84.75 82.02 53.97 52.87 0.544 73.24
+ MS-Mix (ours) 84.76 83.60 84.95 83.47 54.78 53.06 0.543 74.10
ALMT 85.33 81.13 85.31 81.66 54.71 52.84 0.542 73.50
+ Manifold Mix 86.54 83.69 86.37 83.90 53.38 51.51 0.542 74.23
+ MultiMix 85.64 81.73 85.65 82.26 54.04 54.39 0.531 73.95
+ $\mathcal{P}ow$Mix* 85.85 - 85.94 - 54.89 53.29 0.535 -
+ $\mathcal{P}ow$Mix 85.77 81.25 85.42 81.92 54.17 52.89 0.541 73.57
+ MS-Mix (ours) 85.91 83.22 85.78 83.48 55.18 53.60 0.540 74.53
GLoMo 85.00 83.44 84.88 83.65 54.21 52.47 0.543 73.94
+ Manifold Mix 85.94 83.85 85.80 84.05 55.16 53.09 0.546 74.65
+ MultiMix 84.97 83.90 84.87 83.72 54.55 51.82 0.547 73.97
+ $\mathcal{P}ow$Mix 85.41 83.22 85.29 83.47 55.12 53.18 0.541 74.28
+ MS-Mix (ours) 85.84 84.08 85.72 84.13 55.41 53.27 0.533 74.74
Table 4. Results of Various Approaches on The SIMS Dataset. Bold: Best performance. Underline: Second-best performance. $\dagger$}: result reported in (Mao et al., 2022). *: result reported in (Georgiou et al., 2024). "-": result was not reported in the original paper.
MODEL ACC2 ACC3 ACC5 F1-score MAE $\downarrow$ Avg.
@@
+ Manifold Mix 79.21 67.83 40.48 78.92 0.443 66.61
+ MultiMix 78.99 66.08 35.45 78.01 0.443 64.63
+ $\mathcal{P}ow$Mix 79.21 67.40 37.42 79.06 0.423 65.77
+ MS-Mix (ours) 81.18 68.71 40.04 81.14 0.421 67.77
@@
+ Manifold Mix 77.90 65.86 36.98 76.90 0.453 64.41
+ MultiMix 76.59 64.55 36.32 76.70 0.444 63.54
+ $\mathcal{P}ow$Mix 77.86 66.96 39.54 77.33 0.441 65.42
+ MS-Mix (ours) 78.34 65.65 40.61 77.92 0.433 65.63

### v2

@@ 4.1 Datasets
The MOSI [42] dataset is an English benchmark for emotion recognition, consisting of 2,199 opinion segments from 93 YouTube movie reviews. It offers aligned text (transcripts), audio, and visual (facial) data, annotated with sentiment intensity on a continuous scale from -3 (strongly negative) to +3 (strongly positive) by five independent annotators, with final labels calculated as the average of the annotators’ scores. The dataset reflects real-world complexity, including modality incongruity (e.g., sarcasm), and contains about 2.2 hours of video from 89 speakers. Widely used in multimodal fusion research, this dataset comprises pre-extracted features (e.g., acoustic data and facial action units) alongside speaker metadata.
@@
The MOSEI [43] dataset includes 23,453 video segments from more than 1,000 different speakers covering over 250 topics. It is the largest multimodal benchmark for sentiment and emotion analysis. The dataset provides aligned text transcripts, audio signals, and visual frames, all annotated by trained annotators with continuous sentiment intensity scores ranging from -3 to 3 (e.g., MOSI) and six emotion intensities (happiness, sadness, anger, fear, disgust, surprise). It captures real-world communication nuances, including cross-modal conflicts, and supports research in multimodal fusion, emotion recognition, and sarcasm detection. Preprocessed features of MOSI and MOSEI are available through the CMU-Multimodal SDK [47].
@@
The SIMS [44] dataset is a Chinese multimodal benchmark featuring 2,281 video segments from 200 diverse online videos (vlogs, reviews, dialogues). It provides aligned text transcripts, audio, and visual data, with unified sentiment scores on a continuous scale from -1 (strongly negative) to +1 (strongly positive). As the first Chinese dataset to emphasize cross-modal dynamics, SIMS supports research on asynchronous multimodal fusion, cross-modal interaction modeling, and Mandarin sentiment analysis, with raw videos and preprocessed features publicly available.
4.2 Backbones and Baseline Methods
To enable a comprehensive and systematic evaluation, we compared the proposed method against four representative mixup-based augmentation approaches: Manifold Mixup [34] performs random mixing in the latent space to encourage smoother decision boundaries; MultiMix [18] is specifically designed for multimodal fusion through modality-specific mixing operations; $\mathcal{P}$owMix [17] introduces a weight-aware mixing strategy that dynamically adjusts the mixing ratios according to the importance of each modality. AdvMixUp [33] leverages adversarial training to generate sample-dependent hard mixed samples using feature-level masks. We extend this method to the MSA task by adapting the input from raw images to features and applying modality-specific mixing for each modality.
These comparisons were conducted across six widely-used MSA architectures: TFN [21], LMF [22], MuIT [12], MISA [46], ALMT [5], and GLoMo [6]. These backbones encompass a broad range of feature extraction paradigms and fusion strategies, spanning from early influential works to more recent state-of-the-art approaches in MSA. All models were implemented using available code and evaluated under identical settings using the M-SENA framework [4].
4.3 Experimental Settings
@@
We use the evaluation metrics defined in the M-SENA unified framework [4] to ensure consistent, comparable results. These include binary accuracy ($ACC_{2}$) and F1-score (F1), multiclass accuracy ($ACC_{5}$, $ACC_{7}$) and Mean Absolute Error (MAE). Among these, MAE is a regression metric. Classification accuracy values are derived by mapping the regression scores to discrete sentiment categories. The prefix ”w-” denotes the absence of neutral-labeled samples. Furthermore, to facilitate intuitive and comprehensive comparisons across metrics, we have presented the average values (Avg.) for all ACC and F1 metrics separately.
@@
The experiments utilized high-performance servers equipped with an NVIDIA GeForce RTX 3090 Ti or an NVIDIA H20 GPU. The code was implemented in Python 3.10.18 using the PyTorch 2.7.1 framework with CUDA 12.6 support. In MS-Mix, the base mixing ratio $\lambda_{base}$ is sampled from a symmetric Beta distribution with $\alpha=2.0$, and the similarity threshold $\delta$ is fixed at 0.2. The loss weights $\xi_{1}$ and $\xi_{2}$ are set to 0.7 and 0.5, respectively. Additionally, the number of attention heads $h$ is 4 for the MOSI and MOSEI and 6 for the SIMS dataset.
Table 1: Results of Various Approaches on The MOSI Dataset (Mean ± std over 5 Independent Runs).
Bold: Best performance. $\dagger$}: result reported in [4].
[…]
@@
+ Manifold Mix 79.26±1.07 78.13±1.12 79.07±1.04 78.15±1.11 39.02±1.15 35.23±0.74 0.921±0.012 64.81±0.95
+ MultiMix 79.43±0.63 78.22±1.51 79.31±1.05 78.03±1.12 39.48±0.71 34.95±0.26 0.925±0.013 64.90±0.57
+ AdvMixUp 78.48±2.56 78.13±0.38 79.50±1.17 77.74±1.41 39.58±0.70 34.98±0.63 0.923±0.006 64.73±0.51
+ $\mathcal{P}ow$Mix 79.64±1.42 78.18±0.83 79.35±0.96 77.95±0.97 40.11±1.08 36.16±0.44 0.919±0.006 65.23±0.31
+ MS-Mix (ours) 80.22±0.55 78.72±0.55 79.99±1.11 78.39±0.67 40.45±0.68 35.89±0.20 0.914±0.007 65.61±0.29
@@
+ Manifold Mix 79.23±1.12 77.85±1.28 79.34±0.94 77.95±0.68 37.84±1.22 34.17±0.32 0.943±0.009 64.40±0.21
+ MultiMix 79.92±0.81 78.19±0.24 80.29±0.89 78.58±0.46 38.51±0.49 35.80±0.65 0.925±0.006 65.21±0.31
+ AdvMixUp 79.27±1.06 77.84±0.25 79.40±0.62 78.63±0.53 38.16±0.48 34.70±0.90 0.928±0.004 64.67±0.41
+ $\mathcal{P}ow$Mix 80.05±0.45 78.90±0.30 79.86±0.81 78.55±0.54 38.95±0.83 35.91±0.51 0.915±0.009 65.37±0.39
+ MS-Mix (ours) 82.02±0.26 79.16±0.32 82.01±0.16 79.27±0.68 41.90±0.42 36.52±0.50 0.879±0.003 66.81±0.13
@@
+ Manifold Mix 81.01±0.58 80.38±1.26 80.04±1.95 79.34±0.28 41.14±1.67 36.66±1.37 0.864±0.019 66.43±0.83
+ MultiMix 81.49±0.72 80.20±0.67 81.62±0.48 80.19±0.91 42.76±0.87 37.37±0.69 0.856±0.004 67.27±0.47
+ AdvMixUp 81.20±0.55 79.19±0.27 81.64±0.39 79.57±0.74 42.64±0.75 36.51±0.46 0.859±0.007 66.79±0.28
+ $\mathcal{P}ow$Mix 81.44±0.56 79.70±0.62 81.38±0.52 79.53±0.84 41.57±0.62 36.39±0.87 0.876±0.04 66.67±0.27
+ MS-Mix (ours) 82.02±0.28 79.69±0.66 81.97±0.40 79.89±0.84 43.33±0.67 37.48±0.66 0.830±0.006 67.40±0.11
@@
+ Manifold Mix 83.03±0.84 81.49±0.53 83.07±0.84 81.47±0.51 48.00±0.85 41.84±0.38 0.762±0.013 69.82±0.42
+ MultiMix 82.91±0.66 81.80±0.41 82.97±0.91 81.88±0.51 47.40±0.77 42.04±0.36 0.751±0.005 69.83±0.31
+ AdvMixUp 83.44±0.47 81.39±0.50 83.34±0.50 81.34±0.69 47.59±0.59 41.56±0.65 0.762±0.007 69.78±0.26
+ $\mathcal{P}ow$Mix 83.17±0.80 81.45±0.65 83.43±0.46 81.41±0.47 48.16±0.31 42.03±0.55 0.757±0.010 69.94±0.36
+ MS-Mix (ours) 83.94±0.52 82.21±0.63 83.81±0.32 82.14±0.37 48.59±0.56 42.36±0.31 0.746±0.005 70.51±0.19
ALMT 83.48±0.80 82.06±0.61 83.39±0.78 81.94±0.59 51.32±0.75 45.46±0.86 0.733±0.003 71.28±0.23
+ Manifold Mix 84.16±0.76 82.58±0.82 83.58±0.50 82.60±0.77 51.51±0.93 46.04±1.39 0.734±0.008 71.74±0.15
+ MultiMix 83.53±0.71 82.05±0.38 83.31±1.03 81.89±0.61 48.92±0.74 44.21±0.65 0.747±0.005 70.65±0.38
+ AdvMixUp 83.95±0.23 81.94±0.35 83.61±0.67 81.95±0.68 51.44±1.20 46.89±0.59 0.726±0.004 71.63±0.21
+ $\mathcal{P}ow$Mix 83.38±0.50 82.34±0.96 83.26±0.84 82.33±0.93 51.67±0.53 45.72±0.64 0.730±0.007 71.45±0.25
+ MS-Mix (ours) 84.02±0.51 83.21±0.60 83.93±0.33 83.02±0.28 52.33±0.65 46.98±0.37 0.721±0.004 72.25±0.32
GLoMo 83.81±0.27 82.19±0.32 83.73±0.28 82.13±0.33 52.59±0.71 47.10±0.94 0.743±0.010 71.92±0.22
+ Manifold Mix 84.78±0.43 82.95±0.62 84.80±0.23 83.00±0.70 52.89±0.96 49.45±0.99 0.731±0.006 72.98±0.18
+ MultiMix 84.94±0.56 83.23±0.58 85.02±0.49 83.32±0.58 53.01±0.47 50.46±0.93 0.728±0.004 73.33±0.22
+ AdvMixUp 84.27±0.73 83.24±0.65 84.39±0.59 83.25±0.65 53.27±0.49 51.09±0.68 0.730±0.004 73.25±0.36
+ $\mathcal{P}ow$Mix 84.81±0.29 83.36±0.19 84.84±0.36 83.48±0.35 53.14±0.75 49.56±2.06 0.730±0.007 73.20±0.31
+ MS-Mix (ours) 85.21±0.63 83.46±0.19 85.23±0.56 83.43±0.32 54.64±0.96 52.31±0.59 0.726±0.005 74.04±0.47
4.4 Verification Performance of MS-Mix
This chapter evaluates MS-Mix on three publicly available MSA datasets: MOSI [42], MOSEI [43], and SIMS [44]. Experiments were conducted using six backbone models, representing both classical [21, 22, 12, 46] and state-of-the-art approaches [5, 6] in terms of model architectures and fusion strategies, to demonstrate the general applicability of MS-Mix across different frameworks. For each model, we compared against classical Mixup variants, including Manifold Mixup (Manifold Mix) [34] and Multimix [18], as well as a recently proposed mixup-based method in MSA, $\mathcal{P}ow$Mix [17], and the adversarial-based Mixup method, AdvMixUp [33].
To enhance the reliability and robustness of our results, all experiments were conducted five times with different random seeds, and we report the mean and standard deviation (std) for each metric. As shown in Tables 1 to 3, MS-Mix achieves top performance, ranking first or second across all metrics compared with the baseline and other mixup-based methods. It is particularly noteworthy that MS-Mix consistently outperforms all competing approaches on the comprehensive Avg. metric across every backbone model.
Table 2: Results of Various Approaches on the MOSEI Dataset (Mean ± std over 5 Independent Runs).
Bold: Best performance. $\dagger$}: result reported in [4].
@@ 4.4 Verification Performance of MS-Mix
+ Manifold Mix 83.41±0.73 81.42±0.48 83.42±0.61 81.51±0.65 52.87±0.65 51.25±0.72 0.571±0.003 72.31±0.19
+ MultiMix 82.73±0.50 80.61±0.63 82.46±0.64 80.46±0.98 53.17±0.74 51.74±0.64 0.569±0.003 71.86±0.32
+ AdvMixUp 82.85±0.51 80.77±0.71 82.96±0.65 80.98±0.54 53.32±0.63 52.20±0.69 0.566±0.002 72.18±0.16
+ $\mathcal{P}ow$Mix 83.36±1.00 81.25±0.65 83.37±0.72 81.51±0.53 53.08±0.58 51.64±0.55 0.564±0.004 72.37±0.56
+ MS-Mix (ours) 83.41±0.39 82.16±0.26 83.30±0.49 82.07±0.31 53.87±0.44 52.20±0.29 0.559±0.002 72.84±0.17
@@
+ Manifold Mix 83.21±0.64 78.91±1.49 83.39±0.68 79.01±1.46 53.40±0.83 52.04±0.76 0.572±0.008 71.66±0.44
+ MultiMix 84.17±0.71 80.73±0.80 84.29±0.58 80.80±0.83 53.72±0.88 52.72±0.96 0.564±0.002 72.74±0.36
+ AdvMixUp 84.12±0.55 81.72±0.81 84.12±0.63 81.79±0.75 53.47±1.04 52.14±0.77 0.565±0.008 72.90±0.38
+ $\mathcal{P}ow$Mix 83.95±0.31 80.97±0.59 84.02±0.30 81.27±0.53 54.37±0.59 52.73±0.74 0.560±0.006 72.88±0.30
+ MS-Mix (ours) 84.53±0.54 82.99±0.48 84.60±0.40 83.04±0.38 54.45±0.43 52.65±0.47 0.556±0.004 73.71±0.27
@@
+ Manifold Mix 84.39±0.27 81.31±1.19 84.32±0.25 81.55±1.01 53.94±0.48 52.39±0.48 0.562±0.002 72.98±0.58
+ MultiMix 83.99±0.36 81.45±0.68 84.03±0.53 81.75±0.61 54.22±0.52 52.46±0.49 0.565±0.006 72.98±0.19
+ AdvMixUp 84.31±0.50 81.17±0.17 84.29±0.71 80.94±0.74 54.30±0.44 51.98±0.17 0.570±0.003 72.83±0.33
+ $\mathcal{P}ow$Mix 84.71±0.62 81.74±0.75 84.65±0.90 81.85±0.35 54.37±0.51 52.68±0.47 0.559±0.003 73.33±0.23
+ MS-Mix (ours) 84.98±0.34 82.08±0.37 84.87±0.73 82.19±0.53 55.00±0.39 53.04±0.28 0.555±0.003 73.69±0.27
@@
+ Manifold Mix 84.51±0.81 80.24±2.22 84.57±0.73 81.18±1.99 54.49±0.55 52.83±0.69 0.548±0.006 72.90±0.90
+ MultiMix 84.23±0.53 81.31±0.47 84.20±0.68 81.26±0.33 54.07±0.47 52.66±0.38 0.552±0.002 72.95±0.33
+ AdvMixUp 84.65±0.58 81.69±0.49 84.55±0.93 81.64±0.64 54.79±0.57 52.69±0.35 0.544±0.009 73.34±0.19
+ $\mathcal{P}ow$Mix 84.49±0.49 82.03±0.29 84.19±0.38 81.95±0.19 54.79±0.46 52.20±0.62 0.547±0.003 73.27±0.18
+ MS-Mix (ours) 84.84±0.31 83.15±0.36 84.72±0.33 83.08±0.51 55.04±0.34 52.95±0.43 0.542±0.004 73.96±0.28
ALMT 85.15±0.80 81.35±0.56 85.04±0.48 81.53±0.50 54.11±0.29 52.73±0.32 0.550±0.003 73.32±0.14
+ Manifold Mix 85.13±1.46 82.16±1.06 85.08±1.33 81.65±2.73 54.33±0.83 52.65±1.00 0.543±0.005 73.33±1.19
+ MultiMix 85.10±1.02 82.00±0.73 85.22±0.60 81.96±0.78 54.70±0.74 52.70±0.70 0.538±0.007 73.61±0.50
+ AdvMixUp 85.22±0.56 81.99±0.32 85.05±0.64 82.03±0.42 54.46±0.54 52.87±0.30 0.544±0.008 73.60±0.32
+ $\mathcal{P}ow$Mix 85.55±0.46 82.37±0.45 85.48±0.51 82.35±0.26 54.58±0.69 53.13±0.45 0.539±0.004 73.91±0.23
+ MS-Mix (ours) 85.41±0.62 82.90±0.66 85.62±0.55 83.08±0.56 55.17±0.21 53.55±0.36 0.536±0.003 74.29±0.30
GLoMo 84.93±0.80 82.71±0.56 85.28±0.50 82.95±0.77 53.80±0.48 52.67±0.60 0.542±0.004 73.72±0.52
+ Manifold Mix 84.88±0.90 83.65±0.57 85.13±0.72 83.28±0.50 56.08±0.84 53.37±1.53 0.535±0.011 74.40±0.48
+ MultiMix 85.13±0.39 83.29±0.67 85.25±0.39 82.92±0.81 55.05±0.28 52.93±0.33 0.535±0.005 74.09±0.27
+ AdvMixUp 85.24±0.40 83.20±0.39 84.92±0.42 82.98±0.35 55.00±0.39 53.04±0.36 0.537±0.004 74.06±0.22
+ $\mathcal{P}ow$Mix 85.08±0.20 83.19±0.23 85.12±0.26 83.38±0.48 55.08±0.29 53.07±0.27 0.533±0.003 74.15±0.07
+ MS-Mix (ours) 85.32±0.26 84.12±0.31 85.25±0.35 84.17±0.38 55.84±0.22 53.15±0.30 0.531±0.003 74.64±0.10
Table 3: Results of Various Approaches on The SIMS Dataset (Mean ± std over 5 Independent Runs).
Bold: Best performance. $\dagger$}: result reported in [4].
MODEL ACC2(%) $\uparrow$ ACC3(%) $\uparrow$ ACC5(%) $\uparrow$ F1(%) $\uparrow$ MAE $\downarrow$ Avg(%) $\uparrow$.
@@
+ Manifold Mix 78.56±1.13 66.77±0.91 39.54±0.73 78.62±1.01 0.437±0.006 65.87±0.72
+ MultiMix 79.61±0.62 65.95±0.80 37.23±0.86 79.71±0.74 0.438±0.003 65.63±0.26
+ AdvMixUp 79.42±0.96 67.15±1.48 39.33±0.35 79.43±0.62 0.427±0.007 66.33±0.73
+ $\mathcal{P}ow$Mix 79.85±0.56 67.89±0.67 39.50±0.44 79.95±0.52 0.428±0.003 66.80±0.29
+ MS-Mix (ours) 80.66±0.47 68.39±0.72 40.74±0.52 80.71±0.73 0.423±0.005 67.62±0.39
@@
+ Manifold Mix 78.08±0.82 65.85±0.68 38.48±0.92 78.04±0.49 0.447±0.004 65.11±0.52
+ MultiMix 77.97±0.27 65.55±0.64 37.61±1.04 77.70±0.51 0.448±0.011 64.71±0.30
+ AdvMixUp 78.09±0.71 65.96±0.47 39.49±0.89 78.16±0.68 0.435±0.007 65.43±0.63
+ $\mathcal{P}ow$Mix 77.91±0.41 65.88±0.17 39.64±0.49 78.04±0.42 0.437±0.002 65.37±0.25
+ MS-Mix (ours) 78.12±0.24 65.80±0.30 41.02±0.82 78.16±0.37 0.433±0.003 65.77±0.20